OpenStudy (anonymous):

prove tha the locus of the point of interfsection of the line drawn through the points (a,0) , (-a,0) with include a constant angle β is the circle x^2+y^2-a^2±2acotβ

OpenStudy (anonymous):

step1 let the moving point be (h,k)

OpenStudy (anonymous):

step2 so the slope of the line joining (h,k) and (a,0) is (k-0)/(h-a) so the slope of the line joining (h,k) and (-a,0) is (k-0)/(h+a)

OpenStudy (anonymous):

step 3 By the given condition we have +- tanβ = (k/(h-a) - k/(h+a) )/(1 + k^2/(h^2-a^2)..

OpenStudy (anonymous):

step 4 simplify the equation both sides and replace h by x and k by y to get the desired equation

OpenStudy (anonymous):

hence we have +- tanβ = 2ak /(h^2+k^2-a^2) or (h^2+k^2-a^2) +- 2akcotβ =0 or x^2+y^2 -a^2+-2aycotβ =0

OpenStudy (anonymous):

thank you very much but could you explain step 3 please?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

in step3 the angle say β between the lines having slopes m1 and m2 is given by +- tanβ = (m1-m2)/(1+m1*m2)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw